OpenStudy (anonymous):

Okay now this one is more confusing to me. Which of the following is the formula for the geometric sequence? 8, -16, 32, -64,...

OpenStudy (anonymous):

These are the answer choices they have: A. an = 8(-2)n-1 B. an = 2(-8)n-1 C. an = -8(2)n-1 All of the 'an' are supposed to be \[a _{n}\]

OpenStudy (mertsj):

A

OpenStudy (anonymous):

Thanks hun! :)